OpenStudy (anonymous):

PLEASE HELP QUICKLYYY Divide: (6x) / (12y) divided by (4x^5) / (3y^3) so far i flipped them to multiply, but idk where to go from here

OpenStudy (campbell_st):

well the rule for dividing fractions is flip the 2nd fraction and multiply \[\frac{6x}{12y} \times \frac{3y^3}{4x^5}\] which becomes \[\frac{6 \times 3 \times x \times y^3}{12 \times 4 \times x^5 \times y}\] eliminate common factors for the answer

OpenStudy (anonymous):

i have the answer, i just don't understand how to get to it... the answer is (3y^2) / (8x^4) i see how it becomes x^4 and y^2... but why is it 3 and 8?

OpenStudy (campbell_st):

ok... look at the numbers 1st what does 18/48 simplify to...? hint divide them both by 6

OpenStudy (anonymous):

ohhhh. i wasn't multiplying the numbers first, i was dividing.. I understand now, i see what i did wrong.

OpenStudy (anonymous):

Thank you so much!